OpenStudy (anonymous):

HELP! PLEASE! ALREADY ANSWERED! NEED SOMEONE TO CHECK MY ANSWERS!! JUST ONE I NEED HELP WITH! I TRIED TO FIND THE ANSWER, BUT I CAN'T!! --last question on the document.

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

hartnn (hartnn):

first 2 are correct

OpenStudy (hba):

Youv'e done the first one correct.

hartnn (hartnn):

show what u done for 3rd ??

OpenStudy (hba):

You didn't tick the 3rd one ?

OpenStudy (anonymous):

Yep, both are correct!

hartnn (hartnn):

so u got 3 confirmations for 1st 2 :P

OpenStudy (hba):

@charlotteakina Actually you got the first two right.Now just solve the 3rd one and tell me what you get.

OpenStudy (hba):

ok :)

OpenStudy (hba):

How come (-8)^2=-64 ?

OpenStudy (hba):

(-8)^2=-8*-8= (-)(-)=+

OpenStudy (hba):

Yeah :D

OpenStudy (hba):

huh ? Just show me the steps.

OpenStudy (hba):

Continue from here, \[\sqrt{64+36}\]

OpenStudy (hba):

Yay :D

OpenStudy (hba):

Here is your medal :D

OpenStudy (hba):

\[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\] Yeah your'e Actually right :)

hartnn (hartnn):

yes, u r right.